A rolling ball moves from x1 = 3.4 cm to x2 = - 4.2 cm during the time from t1 = 3.0 s to t2 =

6.1 s. What is its average velocity?


1
Expert's answer
2021-10-11T08:55:14-0400

The average velocity

"v_x=\frac{x_2-x_1}{t_2-t_1}=\frac{-4.2-3.4}{6.1-3.0}=-2.5\:\rm m/s"